# Antinucleon-nucleon interactions in covariant chiral effective field theory

### v1

@@ III Scattering equation and phase shifts
We separately solve the Kadyshevsky in the isospin basis for $I=0$ and $I=1$ and fit the resulting phase shifts to those in Ref. Zhou and Timmermans (2012) to determine the corresponding LECs. To remove the ultraviolet divergences, the potential is regularized with a non-local Gaussian-type cut-off function,
@@
with the cut-off value varied in the range $\Lambda=450-600$ MeV. We use the non-local cut-off function so that the contribution of the contact term in each partial wave is not mixed.
@@ IV Results and discussions
Table 1: LO low-energy constants for different cutoffs. The parameters related to the scattering process are in units of $10^{4}$ GeV-2, while the parameters related to the annihilation process are in units of $10^{2}$ GeV-1
@@
The phase shifts obtained in our study, the NLO non-relativistic results Kang et al. (2014), and the $\bar{N}N$ PWA Zhou and Timmermans (2012) for laboratory energies up to $200$ MeV are shown in Figs. 2-5. The partial waves are labeled in the spectral notation ${}^{\left(2I+1\right)\left(2S+1\right)}L_{J}$, and the bands are generated by varying the cutoff in the range $\Lambda=450-600$ MeV for both the relativistic and non-relativistic calculations. The LO non-relativistic phase shifts are not included for comparison because the annihilation potential, in this case, is only non-zero for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves. Hence, the descriptions of the phase shifts of other partial waves are very bad. In addition, even for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves, the differences between the LO non-relativistic phase shifts and the PWA are significant compared with the differences between the NLO non-relativistic phase shifts and the PWA.
@@
The LO relativistic results of the $J=0$ partial waves agree with the PWA for the energy region shown here. Compared with its NLO non-relativistic counterpart, the overall cutoff dependence of the LO relativistic phase shifts is weaker, especially for the real parts of the phase shifts of the $I=1$ partial waves. At the same time, one can observe a sizeable cutoff dependence in the NLO non-relativistic results for energies above $150$ MeV because of the resonance-like behaviors. Since the number of free parameters for the $J=0$ partial waves in the LO relativistic and NLO non-relativistic potentials is identical ($4$ for the ${}^{1}S_{0}$ partial wave and $2$ for the ${}^{3}P_{0}$ partial wave, including annihilation parameters), the relatively weaker cutoff dependence has to do with the relativistic corrections of the scattering equation and the scattering potentials at orders higher than $\mathcal{O}\left(p^{2}\right)$ (in the conventional Weinberg power counting). Note that the NLO non-relativistic results are obtained by fitting to the PWA of Ref. Zhou and Timmermans (2012) at $p_{\text{lab}}\leq 500$ MeV, while in our study one more datum at $p_{\text{lab}}=600$ MeV is also included in the fitting process for the ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves as explained above. However, adapting the same fitting strategy in the NLO non-relativistic framework ruins the descriptions of PWA at $p_{\text{lab}}\leq 500$ MeV. Therefore, the improvement in the cutoff dependence in the relativistic framework cannot be completely attributed to the differences in the fitting procedures. An exception exists in the imaginary part of the phase shift of the ${}^{11}S_{0}$ partial wave, where the cutoff dependence of the LO relativistic results is sizeable at laboratory energies above $150$ MeV. This is related to the description of the ${}^{13}P_{1}$ partial wave, whose PWA yields a negative phase at low energies, which tends to become positive at higher energies. As argued in Ref. Kang et al. (2014), reproducing such phase shifts requires a repulsive potential at large separations of the antinucleon and nucleon but becomes attractive at short distances. Since the scattering potential for ${}^{3}P_{1}$ is controlled by the LECs in the ${}^{1}S_{0}$ partial wave as shown in Eq. II, the description of $\delta_{\text{I}}\left({}^{11}S_{0}\right)$ is influenced by the demand for such an attractive potential. Improvement might be possible at NLO.
@@
For the $J=1$ uncoupled channels, the LO relativistic and NLO non-relativistic results are comparable. The NLO non-relativistic result is better for the imaginary part of the phase shift in the ${}^{11}P_{1}$ partial wave. In comparison, the LO relativistic results are better for the real part of the phase shift in the ${}^{31}P_{1}$ partial wave. Still, the relativistic corrections are not attractive enough to account for the discrepancies between the calculated phase shifts and the PWA. As for other partial waves, both results are comparable, but the cutoff dependence of the LO relativistic results is weaker than that of the NLO non-relativistic results at laboratory energies above 150 MeV, in analogy to the results for the ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves. It should be emphasized that the LO relativistic scattering potentials for the ${}^{1}P_{1}$ and ${}^{3}P_{1}$ partial waves are determined by $S$- wave LECs as shown in Eq. (II). In contrast, the NLO non-relativistic scattering potentials contain as many LECs as the annihilation potentials. Thus, the improvements in the cutoff dependence must originate from the relativistic corrections.
@@
For the $J=1$ coupled channels, the $S$-wave phase shifts are generally well reproduced. The ${}^{3}D_{1}$ phase shift and mixing angle $\epsilon_{1}$ show strong cutoff dependence. However, it is not so surprising since they have no free parameters. The intriguing thing is that the relativistic corrections shift the trends of $\delta_{\text{I}}\left({}^{33}D_{1}\right)$ and $\text{Im}\left(\epsilon_{1}\right)$ to the right direction at laboratory energies above $100$ MeV compared to their non-relativistic counterparts. However, the correction seems too large for the ${}^{33}D_{1}$ partial wave. As a result, the cutoff dependence of $\delta_{\text{I}}\left({}^{33}S_{1}\right)$ becomes large at that energy region.
@@
Figure 2: Real and imaginary parts of the phase shift for the ${}^{1}S_{0}$ and ${}^{3}P_{0}$ partial waves. The gray bands show the LO relativistic chiral EFT results with the cutoff in the range $\Lambda$ = 450–600 MeV. The pink bands show the NLO non-relativistic chiral EFT results of Ref. Kang et al. (2014). The blue dots refer to the solution of the PWA of Ref. Zhou and Timmermans (2012).
@@
Next, we turn to the near-threshold $\bar{N}N$ structures. The phase shifts shown in Figs. 25 suggest the existence of bound states in the ${}^{11}S_{0},^{13}P_{0},^{13}S_{1},$ and ${}^{33}S_{1}$ channels because their phase shifts are about $180^{\circ}$ at threshold. Therefore, we search for possible bound states at the energy region near the $\bar{N}N$ threshold. The corresponding binding energies obtained with our relativistic potential and the NLO non-relativistic potential Kang et al. (2014) are summarized in Table 2. Although these structures have complex $E_{B}$, and the sign of the real part of $E_{B}$ is even positive in some cases, according to Refs. Kang et al. (2014); Badalian et al. (1982), the poles that we found can still be referred to as bound states because they lie on the physical sheet and move below the threshold when the annihilation potential is switched off. Note that compared to the NLO non-relativistic results, a bound state emerges in the ${}^{33}S_{1}$ channel with a relatively large width, which reflects the differences in the potentials of the ${}^{33}S_{1}-^{33}D_{1}$ coupled channel. Moreover, We also find a deeply bound state with $E_{B}=\left(-102.2,-152.5\right)-\text{i}\left(79.1,199.3\right)$ MeV in the ${}^{11}S_{0}$ channel, whose quantum number is consistent with the pseudoscalar interpretation of $X(1835)$, $X(1840)$, and $X(1880)$ suggested by the BESIII Collaboration Ablikim et al. (2005, 2024), despite that it is located far below the $\bar{N}N$ threshold and our result suffer relatively large uncertainties. A firm conclusion can only be drawn once reliable theoretical uncertainties can be estimated. We want to mention that the studies employing the semi-phenomenological $\bar{N}N$ interactions have found a bound state in the ${}^{11}S_{0}$ channel Yan et al. (2005); Sibirtsev et al. (2005); Ding and Yan (2005); Dedonder et al. (2009), although the predicted binding energies are rather different. Therefore, a NLO study is needed to confirm the nature of this state. Apart from the bound states, the phase shifts exhibit resonance-like structures in ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves at energies above $150$ MeV. Thus, we also look for poles in the second Riemann sheet in these two channels. However, we do not find any resonant states in this energy region.
@@
LO relativistic NLO non-relativistic Kang et al. (2014)
${}^{11}S_{0}$ $\left(-102.2,-152.5\right)-\text{i}\left(79.1,199.3\right)$ No near-threshold structure
@@
${}^{33}S_{1}$ $\left(-17.6,7.0\right)-\text{i}\left(128.9,134.4\right)$ No near-threshold structure
@@ V Summary and Outlook
We have studied the $\bar{N}N$ interaction in the covariant chiral effective field theory. The $\bar{N}N$ potential was calculated at LO, and the corresponding LECs were determined by fitting to the phase shifts and inelasticities provided by the PWA of the $\bar{p}p$ scattering data Zhou and Timmermans (2012). The overall description of the PWA with the LO relativistic potential is comparable to that obtained with the NLO non-relativistic potential, similar to the situation observed in the $NN$ interaction. In addition, we searched for near $\bar{N}N$ threshold structures, and found several bound states in the ${}^{11}S_{0},^{13}P_{0},^{13}S_{1}$, and ${}^{33}S_{1}$ channels. The quantum number of ${}^{11}S_{0}$ supports the pseudoscalar interpretation of $X(1835)$, $X(1840)$, and $X(1880)$ observed by the BESIII Collaboration. However, the mass of this bound state is much smaller than $X(1835)$, $X(1840)$, and $X(1880)$. A NLO study is needed to confirm the nature of this state. We note that there is a bound state with a binding energy $E_{B}=\left(-17.6,7.0\right)-\text{i}\left(128.9,134.4\right)\text{MeV}$ in the ${}^{33}S_{1}$ channel, which is, however, missing in the non-relativistic interaction.
Although the $\bar{p}p$ data can be described reasonably well in the relativistic approach, comparable to or even slightly better than the NLO non-relativistic results, further refinements can still be made. For example, the annihilation potential is approximated in the conventional Weinberg power counting, the theoretical uncertainties are estimated roughly by varying the cutoff, and full renormalization group invariance has not been achieved. We will study these issues in the future.
@@ VI Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant No.2023YFA1606700, the National Natural Science Foundation of China under Grants No.12347113, and the Chinese Postdoctoral Science Foundation under Grants No.2022M720360. We thank Dr. Xian-Wei Kang for the enlightening discussions regarding the annihilation potential. Yang Xiao thanks Dr. Chun-Xuan Wang for the valuable discussions.

### v2

@@ III Scattering equation and phase shifts
We separately solve the Kadyshevsky equation in the isospin basis for $I=0$ and $I=1$ and fit the resulting phase shifts to those in Ref. Zhou and Timmermans (2012) to determine the corresponding LECs. To remove the ultraviolet divergences, the potential is regularized with a non-local Gaussian-type cut-off function,
@@
with the cut-off value varied in the range $\Lambda=450-600$ MeV. We want to emphasize that the employed regulator here is sort of “old-fashioned” in the light of the studies on the $NN$ interactions in the conventional ChEFT because it distorts the long-range part of the potential thus slowing down the convergence Reinert et al. (2018). In our framework, although the covariant OPE potential, in contrast to its conventional counterpart, is non-local due to the presence of Dirac spinors, the employed regulator might still introduce some cut-off artifacts because such a regularization implies including a sequence of terms proportional to $\frac{1}{q^{2}-m_{\pi}^{2}}\left(-\frac{p^{6}+p^{\prime 6}}{\Lambda^{6}}\right)^{n}$ in the conventional expansion for nucleons in the unregularized covariant OPE potential. Thus, more careful studies are necessary to understand a proper regularization method in our framework comprehensively. Note that the semi-local regulator has been proven to maintain the long-range part of the pion-exchange potential and remove cut-off artifacts simultaneously, as suggested in Ref. Reinert et al. (2018). However, implementing such a regulator in the present framework is not straightforward because it involves recalculating the covariant pion-exchange potential and projecting it into the $LSJ$ basis. Additionally, such regularization is not expected to change our results significantly, which will be left for future work.
@@ IV Results and discussions
Table 1: Low-energy constants for different cutoffs. The parameters related to the elastic process are in units of $10^{4}$ GeV-2, while the parameters related to the annihilation process are in units of $10^{2}$ GeV-1
@@
The phase shifts obtained in our study, the NLO HB results Kang et al. (2014), and the $\bar{N}N$ PWA Zhou and Timmermans (2012) for laboratory energies up to $200$ MeV are shown in Figs. 2-5. The partial waves are labeled in the spectral notation ${}^{\left(2I+1\right)\left(2S+1\right)}L_{J}$, and the bands are generated by varying the cutoff in the range $\Lambda=450-600$ MeV for both our hybrid method and the HB approach The LO HB phase shifts are not included for comparison because the potential relevant to the annihilation process, in this case, is only non-zero for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves. Hence, the descriptions of the phase shifts of other partial waves are very bad. In addition, even for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves, the differences between the LO HB phase shifts and the PWA are significant compared with the differences between the NLO phase shifts and the PWA.
@@
Our results for the $J=0$ partial waves agree with the PWA for the energy region shown here. Compared with the NLO HB results, the overall cutoff dependence of the hybrid phase shifts is weaker, especially for the real parts of the phase shifts of the $I=1$ partial waves. At the same time, one can observe a sizeable cutoff dependence in the NLO HB results for energies above $150$ MeV because of the resonance-like behaviors. Since the number of free parameters for the $J=0$ partial waves in the hybrid method and the NLO HB formalism is identical ($4$ for the ${}^{1}S_{0}$ partial wave and $2$ for the ${}^{3}P_{0}$ partial wave, including annihilation parameters), the relatively weaker cutoff dependence has to do with the relativistic corrections of the scattering equation and the potentials at orders higher than $\mathcal{O}\left(p^{2}\right)$ (in the conventional Weinberg power counting). Note that the NLO HB results are obtained by fitting to the PWA of Ref. Zhou and Timmermans (2012) at $p_{\text{lab}}\leq 500$ MeV, while in our study one more datum at $p_{\text{lab}}=600$ MeV is also included in the fitting process for the ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves as explained above. However, adapting the same fitting strategy in the NLO HB framework ruins the descriptions of PWA at $p_{\text{lab}}\leq 500$ MeV. Therefore, the improvement in the cutoff dependence in our framework cannot be completely attributed to the differences in the fitting procedures. An exception exists in the imaginary part of the phase shift of the ${}^{11}S_{0}$ partial wave, where the cutoff dependence of the hybrid results is sizeable at laboratory energies above $150$ MeV. This is related to the description of the ${}^{13}P_{1}$ partial wave, whose PWA yields a negative phase at low energies, which tends to become positive at higher energies. As argued in Ref. Kang et al. (2014), reproducing such phase shifts requires a repulsive potential at large separations of the antinucleon and nucleon but becomes attractive at short distances. Since the real parts of the potential for ${}^{3}P_{1}$ are controlled by the LECs in the ${}^{1}S_{0}$ partial wave as shown in Eq. (II), the description of $\delta_{\text{I}}\left({}^{11}S_{0}\right)$ is influenced by the demand for such an attractive potential. Improvement might be possible by using an NLO covariant potential for the elastic process.
@@
For the $J=1$ uncoupled channels, the hybrid and NLO HB results are comparable. The NLO HB results are better for the imaginary part of the phase shift in the ${}^{11}P_{1}$ partial wave. In comparison, the hybrid results are better for the real part of the phase shift in the ${}^{31}P_{1}$ partial wave. Still, the relativistic corrections are not attractive enough to account for the discrepancies between the calculated phase shifts and the PWA. As for other partial waves, both results are comparable, but the cutoff dependence of the hybrid results is weaker than that of the NLO HB results at laboratory energies above 150 MeV, in analogy to the results for the ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves. It should be emphasized that the real parts of the hybrid potentials for the ${}^{1}P_{1}$ and ${}^{3}P_{1}$ partial waves are determined by the $S$- wave LECs as shown in Eq. (II). In contrast, the real parts of the NLO HB potentials contain as many LECs as the imaginary parts. Thus, the improvements in the cutoff dependence must originate from the relativistic corrections.
@@
For the $J=1$ coupled channels, the $S$-wave phase shifts are generally well reproduced. The ${}^{3}D_{1}$ phase shift and mixing angle $\epsilon_{1}$ show strong cutoff dependence. However, it is not so surprising since they have no free parameters. The intriguing thing is that the relativistic corrections shift the trends of $\delta_{\text{I}}\left({}^{33}D_{1}\right)$ and $\text{Im}\left(\epsilon_{1}\right)$ to the right direction at laboratory energies above $100$ MeV compared to the NLO HB results. However, the correction seems too large for the ${}^{33}D_{1}$ partial wave. As a result, the cutoff dependence of $\delta_{\text{I}}\left({}^{33}S_{1}\right)$ becomes large at that energy region.
@@
Figure 2: Real and imaginary parts of the phase shift for the ${}^{1}S_{0}$ and ${}^{3}P_{0}$ partial waves. The gray bands show our results with the cutoff in the range $\Lambda$ = 450–600 MeV. The pink bands show the NLO HB chiral EFT results of Ref. Kang et al. (2014). The blue dots refer to the solution of the PWA of Ref. Zhou and Timmermans (2012).
@@
We want to emphasize that although the interaction constructed in this work is not fully covariant in the sense that the contributions of all the annihilation channels $V_{\bar{N}N\rightarrow X}$ are expanded in powers of antinucleon/nucleon three momenta, the present method is reasonable at least at a qualitative level according to the previous studies in the $NN$ sector Ren et al. (2018). Considering that the contributions irrelevant to the annihilation process are calculated fully covariantly at LO, we think the above discussions on “power counting” and “convergence” are reasonable.
Next, we turn to the near-threshold $\bar{N}N$ structures. The phase shifts shown in Figs. 25 suggest the existence of bound states in the ${}^{11}S_{0},^{13}P_{0},^{13}S_{1},$ and ${}^{33}S_{1}$ channels because their phase shifts are about $180^{\circ}$ at threshold. Therefore, we search for possible $\bar{N}N$ bound states in these channels. The corresponding binding energies obtained with our hybrid potential and the NLO HB potential Kang et al. (2014) are summarized in Table 2. Although these structures have complex $E_{B}$, and the sign of the real part of $E_{B}$ is even positive in some cases, according to Refs. Kang et al. (2014); Badalian et al. (1982), the poles that we found can still be referred to as bound states because they lie on the physical sheet and move below the threshold when the annihilation potential is switched off. Moreover, we find a deeply bound state with $E_{B}=\left(-102.2,-152.5\right)-\text{i}\left(79.1,199.3\right)$ MeV in the ${}^{11}S_{0}$ channel, whose quantum number is consistent with the pseudoscalar interpretation of $X(1835)$, $X(1840)$, and $X(1880)$ suggested by the BESIII Collaboration Ablikim et al. (2005, 2024), despite that it is located far below the $\bar{N}N$ threshold and our result suffer relatively large uncertainties. A firm conclusion can only be drawn once reliable theoretical uncertainties can be estimated. We want to mention that the studies employing the semi-phenomenological $\bar{N}N$ interactions have found a bound state in the ${}^{11}S_{0}$ channel Yan et al. (2005); Sibirtsev et al. (2005); Ding and Yan (2005); Dedonder et al. (2009), although the predicted binding energies are rather different. Therefore, more studies are needed to confirm the nature of this state. Apart from the bound states, the phase shifts exhibit resonance-like structures in ${}^{31}S_{0}$ and ${}^{33}P_{0}$ partial waves at energies above $150$ MeV. Thus, we also look for poles in the second Riemann sheet in these two channels. However, we do not find any resonant states in this energy region.
@@
Hybrid NLO HB  Kang et al. (2014)
${}^{11}S_{0}$ $\left(-102.2,-152.5\right)-\text{i}\left(79.1,199.3\right)$  [^1] It is unclear whether a $\bar{N}N$ bound states can be observed in this channel because the possible structures are only searched for near $\bar{N}N$ threshold.
@@
${}^{33}S_{1}$ $\left(-17.6,7.0\right)-\text{i}\left(128.9,134.4\right)$ 1
@@ V Summary and Outlook
We have studied the $\bar{N}N$ interaction in a hybrid approach, in which the contribution from the elastic process is determined in the LO covariant ChEFT, while the contribution from the annihilation process is described in the NLO HB ChEFT. The corresponding LECs were determined by fitting to the phase shifts and inelasticities provided by the PWA of the $\bar{p}p$ scattering data Zhou and Timmermans (2012). The overall description of the PWA with the hybrid potential is comparable to that obtained with the NLO HB potential. In addition, we searched for near $\bar{N}N$ threshold structures, and found several bound states in the ${}^{11}S_{0},^{13}P_{0},^{13}S_{1}$, and ${}^{33}S_{1}$ channels. The quantum number of ${}^{11}S_{0}$ supports the pseudoscalar interpretation of $X(1835)$, $X(1840)$, and $X(1880)$ observed by the BESIII Collaboration. However, the mass of this bound state is much smaller than $X(1835)$, $X(1840)$, and $X(1880)$. More detailed studies are needed to confirm the nature of this state.
Although the $\bar{p}p$ data can be described reasonably well in the hybrid approach, comparable to the NLO HB results, further refinements can still be made. For example, the potential responsible for the annihilation process is approximated in the conventional Weinberg power counting, the theoretical uncertainties are estimated roughly by varying the cutoff, and full renormalization group invariance has not been achieved. We will study these issues in the future.
@@ VI Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant No.2023YFA1606703, the National Natural Science Foundation of China under Grants No.12435007 and No.12347113, and the Chinese Postdoctoral Science Foundation under Grants No.2022M720360. Yang Xiao and Jun-Xu Lu thank the Fundamental Research Funds for Central Universities for the support. We thank Dr. Xian-Wei Kang for the enlightening discussions regarding the annihilation potential. Yang Xiao thanks Dr. Chun-Xuan Wang for the valuable discussions.